Find the residue of the function f(z)=1/z(ez-1)


1
Expert's answer
2021-09-02T07:25:25-0400

 If "f(z)" has a pole of order "k" at "z=z_0" then


"rez(f,z_0)=\lim\limits_{z\to z_0}\big(\dfrac{1}{(k-1)!}\dfrac{d^{k-1}}{dz^{k-1}}((z-z_0)^kf(z))\big)"

"e^z=1+\dfrac{z}{1!}+\dfrac{z^2}{2!}+\dfrac{z^3}{3!}+..."

"e^z-1=\dfrac{z}{1!}+\dfrac{z^2}{2!}+\dfrac{z^3}{3!}+..."

"\lim\limits_{z\to0}(zf(z))=\lim\limits_{z\to0}(z(\dfrac{1}{z(e^z-1)}))=\lim\limits_{z\to0}(\dfrac{1}{e^z-1})"

"=does\ not \ exist"

"\dfrac{1}{(2-1)!}\dfrac{d^{2-1}}{dz^{2-1}}((z-0)^2f(z))"

"=\dfrac{d}{dz}(z^2(\dfrac{1}{z(e^z-1)}))"

"=\dfrac{d}{dz}(\dfrac{z}{e^z-1})"

"=\dfrac{e^z-1-ze^z}{(e^z-1)^2}"


"rez(f,0)=\lim\limits_{z\to0}\big(\dfrac{e^z-1-ze^z}{(e^z-1)^2}\big)"

"=\lim\limits_{z\to0}\big(\dfrac{e^z-e^z-ze^z}{2e^z(e^z-1)}\big)"

"=\lim\limits_{z\to0}\big(\dfrac{-z}{2(e^z-1)}\big)"

"=\lim\limits_{z\to0}\big(\dfrac{-1}{2e^z}\big)=-\dfrac{1}{2}"


"rez(f,0)=-\dfrac{1}{2}"